### v1

# CryoSAMU: Enhancing 3D Cryo-EM Density Maps of Protein Structures at Intermediate Resolution with Structure-Aware Multimodal U-Nets

Abstract
Enhancing cryogenic electron microscopy (cryo-EM) 3D density maps at intermediate resolution (4-8 Å) is crucial in protein structure determination. Recent advances in deep learning have led to the development of automated approaches for enhancing experimental cryo-EM density maps. Yet, these methods are not optimized for intermediate-resolution maps and rely on map density features alone. To address this, we propose CryoSAMU, a novel method designed to enhance 3D cryo-EM density maps of protein structures using structure-aware multimodal U-Nets and trained on curated intermediate-resolution density maps. We comprehensively evaluate CryoSAMU across various metrics and demonstrate its competitive performance compared to state-of-the-art methods. Notably, CryoSAMU achieves significantly faster processing speed, showing promise for future practical applications. Our code is available at https://github.com/chenwei-zhang/CryoSAMU.
1 Introduction
Cryogenic electron microscopy (Cryo-EM) has become one of the most prevalent techniques in structural biology for determining protein structures, thereby accelerating structure-based drug discovery [1, 2]. Cryo-EM projects a series of 2D images, which are then reconstructed into 3D electron density maps, providing voxelized representations of proteins. While cryo-EM 3D maps serve as the basis for molecular structure determination, using raw maps is usually not possible as they often lack contrast due to various factors, including molecular motion and heterogeneity, imaging artifacts, and incoherent averaging of image data [3]. To address these limitations, various approaches have been developed to enhance map quality by sharpening or modifying map densities [4, 5, 6, 7, 8, 9, 10]. Traditional methods rely on B-factor correction, which can be applied globally [4, 5] and locally [11, 6]. However, these methods struggle with maps exhibiting varying signal-to-noise ratios and lacking prior knowledge (e.g. local resolution) [10].
With recent advancements in deep learning (DL), fully data-driven methods have been developed to automatically enhance raw cryo-EM maps for protein structure modeling. Leveraging neural networks such as convolutional neural networks (CNNs) [12], generative adversarial networks (GANs) [13], and Transformers [14], these methods achieved promising results in map enhancement. Yet, they are not optimized for intermediate-resolution maps (i.e., 4-8 Å [10]) and rely solely on a single modality—the density map itself—during neural network training, overlooking other relevant modalities such as structural information. This limitation restricts their ability to generalize across diverse protein structures and prevents them from fully leveraging complementary biological information. To address these shortcomings, we thus introduce CryoSAMU, a novel approach that combines 3D map features with structural embeddings derived from the pretrained protein language model ESM-IF1 [15] to enhance 3D Cryo-EM density maps with Structure-Aware Multimodal U-Nets.
Our main contributions are:
  • We propose the first multimodal network that integrates structural information into a 3D U-Net model using cross-attention mechanisms for cryo-EM map enhancement.
  • We develop a self-attention-based post-processing procedure for ESM-IF1’s structural embeddings, effectively preserving both chain and residue relationships while maintaining structural integrity.
  • We train CryoSAMU on a curated dataset of joint density maps at intermediate resolution and associated protein structures, optimizing it for map enhancement.
  • We benchmark CryoSAMU against state-of-the-art approaches across various evaluation metrics over diverse tested samples. We achieve competitive level of performance but with significantly faster processing speeds (approximately 4.2 to 16.7 times), making our method well-suited for large-scale and practical applications.
  • Our ablation study demonstrates significant improvement brought from integrating structural information.
2 Related Work
2.1 Existing map enhancement methods
Conventional map enhancement (sharpening) approaches, including Phenix Autosharpen [4] and RELION postprocessing [5], are based on global B-factor correction. This technique enhances the amplitude of high-frequency Fourier components in raw cryo-EM maps. However, global B-factor-based methods encounter difficulties with maps exhibiting heterogeneous local resolutions, often leading to over- or under-sharpening in specific regions. Despite local B-factor-based sharpening algorithms [11, 6] have been developed to alleviate this limitation, these methods still suffer from poor accuracy in estimating the local resolution of maps, which is crucial for precise local B-factor sharpening.
DeepEMhancer [7] is a pioneering DL-based fully automatic method that leverages a 3D U-Net model to mimic local sharpening effects and enhance map features. Subsequently, CryoFEM [8] that employs convolutional neural networks (CNNs) and EM-GAN [9] that utilizes generative adversarial networks (GANs) have been introduced to further enhance cryo-EM maps. Most recently, with the emergence of vision transformers, EMReady [10], which adopts a Swin transformer architecture [16], has shown superior performance in enhancing map quality for accurate protein structure modeling.
2.2 Protein large language models
The advancement of protein large language models (pLLMs) has enabled unprecedented insights into protein structure, function, and evolution [17, 18, 19, 15, 20, 21, 22]. In analogy to human texts, protein sequences are treated as “biological texts” and input into pLLMs to capture contextual information inherent in the sequences. Notable examples of such models include the ESM family [17, 18, 19], which are pretrained on vast datasets of protein sequences using the masked language modeling strategy, allowing them to develop rich representations that encapsulate evolutionary information.
Addressing the inverse problem of predicting protein sequences from given structures, ESM-IF1 [15] has been developed. Trained on 12 million protein structures derived from AlphaFold2 [23], ESM-IF1 predicts protein sequences from backbone atom coordinates. It is specifically designed to encode both sequence and structural information, including backbone geometry, side chain orientations, and secondary structure elements. These traits make ESM-IF1 a compelling choice for generating structure-aware embeddings that complement the map-only modality.
Figure 1: Overview of the CryoSAMU framework. a Generating protein multimodal representations: structure features are derived from a frozen pretrained ESM-IF1 model with self-attention weighting for a fixed-size representation; map voxel features are simulated via resolution-lowering point spread function and partitioned into smaller cubes. b The CryoSAMU architecture. The experimental map is partitioned into smaller cubes and processed by a U-Net with residual blocks and linear attention modules. Structural embeddings are integrated into the bottleneck layer with cross-attention mechanism. The output cubes are reconstructed into the full-size enhanced map.
3 Method
3.1 Dataset of protein structures and density maps
Our dataset was built with a set of cryo-EM density maps at resolutions from 4.0 Å to 7.9 Å from the EMDB databank [24] and their associated protein structures from the PDB databank [25]. To ensure that density maps are properly aligned with their corresponding PDB structures, we excluded maps and PDBs from the dataset if: (i) maps contain extensive regions without or misaligned corresponding PDB structures; (ii) maps contain other macromolecules except proteins; (iii) PDB structures contain backbone atoms only and/or unknown residues. Furthermore, to enhance training efficiency, we measured the correlation between map-PDB pairs using ChimeraX [26], and removed pairs with correlation score lower than 0.65. To avoid data redundancy, we measured the sequence identity between PDB structures, and retained only one if identity is greater than 30 %. As a result, a total of 384 pairs of cryo-EM maps and associated PDB structures remained. Among these data, 247 ($\sim$65 %), 62 ($\sim$15 %), and 75 ($\sim$20 %) map-PDB pairs were selected as training, validation, and test sets, respectively. Details are listed in Supplementary Tables 1-3.
3.2 Multimodal representations of protein structures
Generating 3D target maps from protein structures   For input experimental maps (denoted as ExpMaps) in training and validation sets, we simulated the corresponding target maps (denoted as TgtMaps) from associated protein structures using the StructureBlurrer package in TEMPy2 [27]. The simulation was performed with a grid interval of 1 Å and a resolution cutoff at 2 Å, based on the convolution of atom points with resolution-lowering point spread functions. Given a PDB structure with $M$ atoms, the simulated density $\rho$ at grid point $x$ is calculated by:
$\rho(\mathbf{x})=\sum_{i=1}^{M}\theta Z_{i}e^{-k|\mathbf{x}-\mathbf{r}_{i}|^{2}},$ (1)
where $Z_{i}$ and $\mathbf{r}_{i}$ refer to the atomic number and the position vector of the i-th heavy atom, respectively. Here, $\theta$ is a scaling factor and $k$ is defined based on resolution [10, 28].
Resampling 3D maps   We first resampled both ExpMaps and TgtMaps to 1 Å/voxel since the cryo-EM maps vary in voxel size. Subsequently, we normalized the density values to a range of 0 to 1 using the 99.9th percentile density value of each map. Due to GPU memory constraints, we partitioned ExpMaps and TgtMaps into smaller 3D subvolume pairs (denoted as exp.3D-images and tgt.3D-images) with size of $64\times 64\times 64$, the largest feasible size that allows for a sufficient batch size (See Figure 1a.). To mitigate boundary artifacts during truncation, we applied zero-padding of 64 voxels on each side along all dimensions. As a result, a total of 29829 exp-tgt image pairs were yielded for network training and 4642 for validation.
[…]
Fixed-size representation with attention weighting   Following the generation of embeddings, we implemented self-attention weighting to create fixed-size representations while preserving the intrinsic relationships between chains and residues. To this end, we computed attention weights based on embedding similarity to identify the most informative regions. Specifically, given a PDB structure containing $C$ chains and $R$ residues per chain, its structural embedding derived from ESM-IF1 is denoted as $\mathbf{E}\in\mathbb{R}^{C\times R\times d}$, where $d=512$ is the embedding dimension. We carried out the refinement process in several steps. First, we computed chain-level embeddings by averaging across residues:
$\mathbf{E}_{\text{chain}}=\frac{1}{R}\sum_{j=1}^{R}\mathbf{E}_{:,j,:},\quad\mathbf{E}_{\text{chain}}\in\mathbb{R}^{C\times d}.$ (2)
Next, we computed the similarity matrix to determine the relative importance of each chain:
$\mathbf{S}=\mathbf{E}_{\text{chain}}\cdot\mathbf{E}_{\text{chain}}^{T},\quad\mathbf{S}\in\mathbb{R}^{C\times C},$ (3)
where each element $\mathbf{S}_{ij}$ represents the similarity between chain $i$ and chain $j$: $\mathbf{S}_{ij}=\mathbf{E}_{\text{chain},i}\cdot\mathbf{E}_{\text{chain},j}^{T}$. To derive attention weights, we applied a column-wise softmax function to $\mathbf{S}$:
$\mathbf{W}_{ij}=\frac{\exp(\mathbf{S}_{ij})}{\sum_{k=1}^{C}\exp(\mathbf{S}_{ik})}.$ (4)
We then aggregated these weights across chains to assign a single importance weight per chain:
$\mathbf{w}_{i}=\frac{1}{C}\sum_{j}^{C}\mathbf{W}_{ij},\quad i=1,2,\dots,C.$ (5)
These weights $\mathbf{w}=[\mathbf{w}_{1},\mathbf{w}_{2},..,\mathbf{w}_{C}]$ reflect the relative importance of each chain, and we leveraged them to aggregate chain-level embeddings into a unified representation:
$\mathbf{E}_{\text{pooled}}=\sum_{i=1}^{C}w_{i}\mathbf{E}_{i,:,:},\quad\mathbf{E}_{\text{pooled}}\in\mathbb{R}^{R\times d}.$ (6)
We further measured the importance of each residue in $\mathbf{E}_{\text{pooled}}$ using a residue-level similarity matrix. Following the same procedure as the chain-level weighting, we obtained a scalar weight $\alpha_{j}$ for each residue $j$, where $j=1,2,\dots,R$. Finally, we applied min-max normalization and resampled the embedding $\mathbf{E}_{\text{pooled}}$ based on the attention weights to a fixed-size representation, $\mathbf{E}_{\text{final}}\in\mathbb{R}^{L\times d}$, where $L=800$. When the input length $R>L$, we selected the top-$L$ residues with the highest attention weights. Conversely, when $R<L$, we sorted the residues by their attention weights and repeated them $\lceil L/R\rceil$ times to reach the target length, ensuring each resulting embedding maintains rich representations and consistent dimensions.
3.3 The model architecture
We proposed a structure-aware multimodal 3D U-Net, as depicted in Figure 1b. The network contains an encoder, bottleneck, and decoder, interconnected by skip connections.
Encoder   The input to the encoder is a 3D volume with a single channel. The encoder comprises four hierarchical layers. The first three layers each consist of two residual blocks, with each block incorporating a group normalization, a SiLU activation [29], and a dropout (p=0.2), followed immediately by a linear self-attention module with 4 heads [30] to capture long-range (global) dependencies across voxels. The channel depth progressively increases as features are abstracted. In the fourth layer, only residual blocks are employed, producing a higher-level feature representation without the addition of attention modules.
Bottleneck   At the bottleneck layer, the feature representation is first refined by a residue block and then by a linear self-attention module. Subsequently, a cross-attention block is introduced to fuse and align the volumetric features with structural embeddings using multi-head attention with 4 heads, where queries are derived from the volume features and keys/values from structural embeddings. This process enables structural conditioning while preserving spatial relationships. A second residual block is then applied to further fuse the combined features from both modalities.
Decoder   The decoder follows a symmetric architecture to the encoder. Feature maps are progressively upsampled using nearest-neighbor interpolation combined with 3D convolutions, and skip connections incorporate corresponding features from the encoder. Finally, a group normalization, a SiLU activation, and a concluding 3D convolution project the processed features to a single output channel.
3.4 Network training and inference
Protein structural embeddings provide an additional modality containing structure information, serving as key-value pairs in the attention mechanism when training. However, since these embeddings are unavailable during validation and inference, we implemented a specialized mode in which the network bypasses the cross-attention operation. In this mode, the network relies exclusively on feedforward transformations with residual connections. This design maintains consistency between training and validation/inference phases while preserving the learned feature representations.
Training   During training, CryoSAMU accepts an exp.3D-image and its corresponding structural embedding as input and generates an enhanced 3D image (denoted as pred.3D-image). Previous studies have shown that L1 loss performs well in similar tasks [7, 8]. However, to improve training stability in the presence of noisy data and outliers that are common in cryo-EM maps, we employed the smooth L1 loss to encourage the generator to minimize the difference between the output pred.3D-image, $X$, and the target tgt.3D-image, $Y$:
[…]
Moreover, to enhance the network’s robustness, we employed TorchIO [31] for data augmentation, including random Gaussian noise, anisotropy, and blurring.
Inference   During inference, the input experimental map was first zero-padded and divided into smaller cubes ($64\times 64\times 64$), following the same strategy used for training data. Each cube was then individually processed by the trained neural network to generate the enhanced cube. These enhanced cubes were subsequently reassembled to reconstruct the map as its original dimensions. To prevent loss of spatial information and ensure smooth transition between cubes, only the central $50\times 50\times 50$ voxels from each enhanced cube were used in the final reconstruction, following the method proposed by Si et al. [32].
Implementation   The network was implemented in PyTorch 2.6.0 with CUDA 12.4, running under Python 3.12.8. Training was conducted using a distributed data parallel (DDP) strategy across two computational nodes connected via NVLink, with each node equipped with four NVIDIA A100 GPUs of 40 GB VRAM. This setup supported a maximum batch size of 18 per GPU. The network was trained over 95 epochs, requiring approximately 63 computational hours. The AdamW [33] optimizer was used with an initial learning rate of 0.0001, along with a cosine annealing learning rate scheduler. To improve training performance while maintaining accuracy, automatic mixed precision training was applied. Additionally, gradient clipping (set to 0.5) was applied to prevent gradient explosion.
4 Experiments and Results
Figure 2: Visual and quantitative comparison of deposited (blue) and CryoSAMU-enhanced (green) maps, with superimposed corresponding PDB structures (brown). a, b: Maps are shown at two contour levels. Left: recommended contour level (volume = 85.74e3). Right: higher contour level (volume = 22.57e3). c: RSCC comparisons between deposited and CryoSAMU-enhanced maps. The example protein is a CX3CL1-US28-G11iN18-scFv16 in TL-state (PDB-7RKF, EMDB-24496, reported resolution of 4.00 Å) [34].
We conducted a comprehensive study to assess the performance of CryoSAMU using a test set of 75 intermediate-resolution cryo-EM density maps and associated PDB structures across a wide range of evaluation metrics.
4.1 Visualization and quantification of map enhancement
We first visualized a CryoSAMU-enhanced map alongside its associated deposited map using UCSF ChimeraX [26]. For a fair comparison, both sets of maps were illustrated with the same volume, which requires contour level adjustments owing to differences in their volume ranges. Specifically, we first presented the deposited map at its recommended contour level and volume, then adjusted the contour level of the corresponding CryoSAMU-enhanced map to match the recommended volume. In addition, we also visualized both maps at a higher contour level with the same volume.
As displayed in Figure 2b, CryoSAMU significantly suppressed noise in the lip nanodisc regions (highlighted by dashed boxes in Figure 2a) of the deposited map for EMDB-24496 (PDB-7RKF). Moreover, the deposited map at a smaller volumes missed certain structural regions corresponding to the protein structures, as highlighted by black boxes in Figure 2a. In contrast, the CryoSAMU-enhanced maps exhibited better alignment with the corresponding protein structures, revealing more structural details, as demonstrated by black boxes in Figure 2b. Similar visual results were observed for another protein structure (see Supplementary Figure 1). Furthermore, residue-level real-space correlation coefficient (RSCC) measurements [35] in Figure 2c suggested significant improvements. Specifically, Chains A, B, and C exhibit RSCC increases compared to the deposited map, with correlations rising from 0.835 to 0.860, 0.848 to 0.865, and 0.801 to 0.836, respectively. In addition, 84.9%, 73.5%, and 90.6% of resiudes in Chains A, B, and C, respectively, showcased higher RSCC scores. Consistent RSCC improvements were also observed in other samples (see Supplementary Figure 2).
4.2 Benchmark I: improvement of real and Fourier space correlations
Metric Deposit Autosharpen DeepEMhancer EMReady CryoSAMU(ours)
CryoSAMU
(w/o struct.)
CC_box 0.731 0.679 0.618 0.862 0.834 0.751
CC_peaks 0.750 0.722 0.611 0.774 0.753 0.698
CC_volume 0.594 0.542 0.534 0.729 0.691 0.571
FSC05 6.124 6.147 5.283 4.668 5.108 6.434
Table 1: Comparison of different methods across various metrics. See Section 4.2.
Figure 3: The violin plots for comparison of different methods across four evaluation metrics (see Section 4.2) over 75 test samples.
We then benchmarked CryoSAMU against other state-of-the-art methods, including Autosharpen [4], DeepEMhancer [7], and EMReady [10], in terms of both real-space and reciprocal-space (i.e., Fourier-space) correlations, across a test set of 75 primary maps. For real-space correlation, we computed three correlation metrics using phenix.map_model_cc [35] for each map-model pair (where the model refers to a protein structure): CC_box, CC_volume, and CC_peaks. These metrics differ based on the choice of map regions used in the calculations. CC_box considers the entire map. CC_volume and CC_peaks focus on regions with the highest density values. However, CC_volume selects grid points only around atomic centers, while CC_peaks selects points located anywhere within the volume. For all three metrics, higher values indicate better map performance. For Fourier-space correlation, we computed Fourier shell correlation (FSC) using phenix.mtriage [35], and reported the unmasked map-model FSC05 values. FSC values are typically represented as a function of the inverse map resolution, where lower value indicates better map resolution.
The average real-space CC and FSC values are listed in Table 1. According to the violin plots shown in Figure 3, CryoSAMU-enhanced maps demonstrated significant improvements over the deposited maps in terms of CC_box and CC_volume, with average values increasing from 0.731 to 0.834 and from 0.594 to 0.691, respectively. The average CC_peaks score showed a slightly increase from 0.750 to 0.753. These results indicate that CryoSAMU effectively enhances deposited maps in both the entire region and the highest-density regions. In contrast, maps processed by Autosharpen and DeepEMhancer exhibited lower scores across all three metrics. EMReady showed slightly better improvements than CryoSAMU across all three metrics. For FSC05 scores, CryoSAMU outperformed the deposited map, Autosharpen, and DeepEMhancer, achieving an average value of 5.108 Å. However, it slightly underperformed compared to EMReady, which achieved an average value of 4.668 Å. These results demonstrate that both CryoSAMU and EMReady consistently enhance the deposited maps in terms of correlations in both real and Fourier spaces.
4.3 Benchmark II: improvement of protein structure modeling
As the goal of enhancing cryo-EM density maps is to improve the performance of protein structure modeling from density maps (i.e., map interpretability), we benchmarked protein structures constructed from CryoSAMU-enhanced maps against those processed by other methods. Specifically, we used a standard structure modeling tool, known as phenix.map_to_model [36], to construct protein structures from 20 maps enhanced by the different tested methods. These maps were randomly selected from the test dataset to ensure that they were not exposed during training, as listed in Supplementary Table 4. To evaluate these structures, we used phenix.chain_comparison [36] to compare the constructed structures against their corresponding ground-truth PDB protein structures. We reported two metrics: residue coverage and sequence match. The residue coverage indicates the fraction of residues in the query structure that match the corresponding residues in the target structure within 3.0 Å, regardless of residue type. The sequence match indicates the percentage of matched residues that share identical residue types between the query and target structures.
Method
Residue
Coverage (%)
Sequence
Match (%)
Deposit 31.71 8.42
Autosharpen 16.00 8.13
DeepEMhancer 24.31 10.0
EMReady 31.61 11.38
CryoSAMU(ours) 38.03 9.33
CryoSAMU(w/o struct.) 8.08 8.13
Table 2: Comparison of average residue coverage and sequence match across different methods.
The average metric scores from all methods are listed in Table 2. Figure 4a and b provide a detailed comparison for each individual test examples in terms of residue coverage and sequence match, respectively. The polar plots clearly showcase that after CryoSAMU enhancement, 19 out of 20 samples exhibited an improvement in residue coverage on deposited maps, with the average score increasing from 31.71% to 38.03%; and 55% of samples exhibited an improvement in sequence match, with the average score increasing from 8.42% to 9.33%. Furthermore, we benchmarked CryoSAMU against other methods, as shown in Figure 4c and d. CryoSAMU achieved the highest residue coverage score among all methods, although its correlation scores were slightly lower than those of EMReady. The sequence match score of CryoSAMU was slightly lower than EMReady and DeepEMHancer, while still better than the deposited maps. These results demonstrate that CryoSAMU enhancement boosts protein structure modeling performance.
Figure 4: a-b: The polar plots for comparison of protein structures constructed from deposited (blue) and CryoSAMU-enhanced (green) maps, using metrics of (a) residue coverage and (b) sequence match. c-d: The box-whisker plots for comparison of different methods across two evaluation metrics over 20 test samples. See Section 4.3.
4.4 Benchmark III: processing time
To evaluate the scalability of CryoSAMU in practice, we recorded the time required to generate each enhanced map of all 75 test samples and compared it against the processing time of other methods. Figure 5 shows the wall-clock time plotted against the volume size of input experimental maps, ranging from the order of $10^{6}$ to $10^{8}$$^{3}$. For a fair comparison, all methods were run on the same workstation equipped with an AMD Ryzen Threadripper 2950X Processor of 32 CPUs and an NVIDIA GeForce RTX 2080 Ti of 12 GB VRAM. Each method was executed with the maximum batch size that our GPU can accommodate: approximately 12 for DeepEMhancer, 64 for EMReady, and 24 for CryoSAMU.
Figure 5: The scatter plot of map processing time against map volume. Each dot represents the processing time for an individual map based on its volume. The shaded area around the regression line denotes the confidence interval of the regression estimate.
CryoSAMU (shown in green) displayed the minimum processing time across maps of varying volumes. Its weak linear dependency on map volume and tight confidence interval around its fit line indicate that CryoSAMU has both optimal scalability and consistent performance. In contrast, DeepEMhancer exhibited a strong linear correlation between processing time and map volume, indicating poor scalability as volume size increases. EMReady showed a wider confidence interval in its linear fit, reflecting high variability in processing time. Notably, several outliers at lower volumes showed significantly longer processing time compared to other methods. For a significantly large map with a volume size of $1.25\times 10^{8}$$^{3}$, CryoSAMU took only 116.48 seconds for generating an enhanced map, while Autosharpen, DeepEMHancer, and EMReady took 552.19, 2963.10, and 1731.719 seconds, respectively. Table 3 lists the average processing time for each method. CryoSAMU achieved an average processing time of 32.49 seconds, approximately 13.6 times faster than EMReady, while generating comparably enhanced maps. These results suggest that CryoSAMU scales efficiently with increasing map volume, making it a promising tool for practical applications.
Autosharpen DeepEMhancer EMReady CryoSAMU
138$\pm$118 544$\pm$517 441$\pm$500 32$\pm$23
Table 3: Average processing time in seconds of different methods.
4.5 Ablation study
We finally conducted an ablation study to evaluate the impact of integrating structural modality. We compared CryoSAMU with (w/) and without (w/o) structural embeddings using 75 test samples for correlation evaluation and 20 test samples for protein structure modeling assessment.
Figure 6 shows that CryoSAMU (w/) outperforms CryoSAMU (w/o) in both CC_box and FSC05, with improvements of 98.7% and 77.3%, respectively, as indicated by scatter points above the diagonal line. Moreover, the most significant gains (points far from 1.0) were observed in poorer-quality deposited maps (colored in purple), which tend to have lower deposited CC_box or higher FSC05 values. Table 1 lists the average real- and Fourier-space metrics, indicating that incorporating structural embeddings derived from ESM-IF1 led to a significant improvement of map enhancement, as also reported in Figure 3. In terms of protein structure modeling, residue coverage significantly raised from 8.08% to 38.03%, while sequence match raised from 8.13% to 9.33% with the integration of structural embeddings. This suggests that structural information helps complement map regions with poor resolutions, artifacts, or noise, thereby increasing the completeness (higher residue coverage) and improving accuracy (higher sequence match) during structure modeling. These findings underscore the importance of integrating structural modality to enable the network to develop structural awareness beyond learning solely from 3D density maps.
Figure 6: Pairwise comparison of enhanced/deposited ratios for CryoSAMU (w/) and (w/o). Each point represents a single map. Point colors encode the deposited map quality.
5 Conclusion and Discussion
In this work, we introduce CryoSAMU, the first structure-aware multimodal network for enhancing cryo-EM density maps at intermediate resolution of protein structures. Our approach combines 3D map features with corresponding structural features through cross-attention mechanisms. In addition, we develop a self-attention weighting algorithm to produce fixed-size representations of structural embeddings derived from the pretrained ESM-IF1 model, preserving inter-chain and residue relationships while maintaining structural integrity. Our benchmark results demonstrate that CryoSAMU preforms competitively with existing cutting-edge methods, closely approaching the performance of EMReady, the current leading tool for cryo-EM density map enhancement. Notably, CryoSAMU achieves the fastest processing speed among all tested methods, positioning it as a promising solution for large-scale and practical applications. Furthermore, our ablation study reveals that incorporating an additional structural modality significantly boosts CryoSAMU’s performance across all evaluation metrics, suggesting a new avenue for future cryo-EM research to explore the effective integration of multimodal data during network training.
Despite CryoSAMU demonstrating superior performance in enhancing cryo-EM maps, its current architecture—based on residual convolutions within a U-Net framework—is primarily designed to capture local information. In practice, capturing global context and long-range dependencies across map voxels could further improve performance. This could be addressed by adopting more hierarchical architectures, such as the Swin Transformer [16], which facilitates feature extraction over larger receptive fields. Moreover, incorporating supplementary loss terms, such as the Structural Similarity Index Measure (SSIM) loss, could mitigate overfitting and enhance training efficiency [10]. These will be explored in our future work. Furthermore, we plan to expand our dataset by including high-resolution maps, which could increase the robustness of the model and further elevate its performance.
Acknowledgments
This work is supported in parts by a NFRFE-2019-00486 grant.
Supplementary Material
Figure 1: Visualizations of deposited (blue) and CryoSAMU-enhanced (green) maps. The corresponding PDB structures (brown) are superimposed on the maps. a Human Dispatched-1 (PDB-6XE6, EMDB-22144, reported resolution of 4.53 Å) [37]. a-b: Maps displayed at the recommended contour level. c-d: Maps displayed at a higher contour level. Visualizations were produced by UCSF ChimeraX [26]. The protein structure modeling completeness and accuracy improved after CryoSAMU enhancement. For instance, residue coverage increased from 53.5% to 64.5%, as well as sequence match increased from 6.1% to 7.7%.
Figure 2: The real-space correlation coefficient (RSCC) comparison between deposited and CryoSAMU-enhanced maps. Top: a transmembrane nanopore TMH4C4 (PDB-6M6Z, EMDB-30126, reported resolution of 5.9 Å) [38]. Bottom: a yeast cytoplasmic exosome (PDB-5G06, EMDB-3366, reported resolution of 4.2 Å) [39]. In the first example of PDB-6M6Z, both Chain B and Chain D exhibited significant RSCC improvements compared to the deposited maps, increasing from 0.706 to 0.768 and from 0.739 to 0.777, respectively. In addition, 85.2% of residues in Chain B and 70.9% of residues in Chain D showed an increase in their RSCC scores. In the second example of PDB-5G06, the average RSCC scores increased from 0.601 to 0.635 for Chain B, with 82.9% of 240 residues showing improvement; and increased from 0.581 to 0.615 for Chain E, with 76.2% of 265 residues perform better.
Table A1: List of all EMDB/PDB examples in training sets.
Table A2: List of all EMDB/PDB examples in validation sets.
Table A3: List of all EMDB/PDB examples in test sets.
Table A4: List of all EMDB/PDB examples for protein structure modeling.